# Prediction intervals for overdispersed binomial endpoints and their application to toxicological historical control data

### v3

@@ 1 Introduction
Despite the fact, that the HCL based validation of a current control is required by several test guidelines (e.g. OECD 471, OECD 473, OECD 487, OECD 490), most guidelines fail to provide reproducible methodology on how to compute HCL. Hence, many heuristics such as the historical range (the min and the max of the HCD) or the mean $\pm$ 2 standard deviations are used for the practical calculation of HCL, but suffer from severe theoretical drawbacks (Menssen 2023). However, this problem can be overcome by the application of prediction intervals, which, based on historical observations (the HCD) directly aim to predict one (or more) value(s) which are derived from the same data generating process (the current control group). Hence several authors recommend the use of prediction intervals for the application of HCL (EFSA 2024, Menssen and Schaarschmidt 2019, Menssen and Schaarschmidt 2022, Menssen 2023, Kluxen et al. 2021, Dertinger et al. 2023).
However, for the proper application of prediction intervals, it is crucial to account for the hierarchical design of the historical control data: Due to the fact that certain experimental units belong to a certain control group (blue panels in fig 1), HCD is comprised of at least one level of hierarchy. Sometimes it might happen, that further levels of hierarchy are present within each control (or treatment) group e.g. that certain animals within a control group belong to a certain cage (shape of the dots in 1). Since each trial is comprised of different living experimental units (cells or animals) standardization between the trials (with respect to the experimental units) is only possible up to a certain extent. Furthermore, lab personnel, housing conditions or animal (or cell) suppliers might change between different historical studies.
Figure 1: Scheme of the hierarchical design of toxicological HCD and the current trial. Blue panels: Control groups of different toxicological studies. Grey panels: Treatment groups of the different toxicological studies, which are usually not taken into account for HCD applications. Black dots: Individual experimental units. Shape of the dots: Possible within group randomization.
The fact that different experimental units belong to different control groups gives rise to systematic between study variation, which usually reflects both: Controllable causes (e.g. differences in handling of experimental units) as well as uncontrollable causes (e.g. the biological variation within the population of the experimental units). This means, that the sheer presence of systematic between study variation is not problematic, but it is rather its magnitude that has to be considered (Deringer et al. 2023). However, this matter needs assay specific discussions by the toxicological community which are off-topic of this manuscript.
@@
It has to be noted that the prediction intervals presented in several standard textbooks and other scientific publications (e.g. Hahn and Meeker 1991, Hahn et al. 2017, Wang 2008, Tian et al. 2022, Nelson 1982) ground on the assumption that the historical sample is unstructured and hence comprised of independent and identically distributed observations, meaning that between-study variability is assumed to be absent. As explained above, this assumption does not make sense in the context of toxicological HCD for which systematic between study variation is a fundamental part of the data generating process, and hence, is regularly reported in real life data. For example, HC data sets that are prone to systematic between study variation can be found in Tarone 1982, Carlus et al. 2013, Menssen et al. 2024, Levy et al. 2019, Tug et al. 2024, Dertinger et al. 2023 (for further details see section 4 of the supplementary material).
@@
In toxicology several study types assess dichotomous endpoints: e.g. Carcinogenicity studies (rats with vs. rats without a tumor), the micro nucleus test (cells with vs. cells without a micronucleus) or the recently developed liquid microplate fluctuation Ames MPFTM test following Spiliotopoulos et al. 2024 (numbers of wells with vs. number of wells without a color change).
@@
It seems natural to model such data based on the binomial distribution. But, as described before, the clustered nature of the HCD gives rise for systematic between-study variation. With other words, if systematic between-study variation is present, the variability of the data will exceed the variance that can be modeled by the simple binomial distribution. This effect is called overdispersion (or extra-binomial variation) and can either be modeled based on the quasi-binomial assumption - which relies on the assumption that the between study overdispersion is constant - or based on the beta-binomial distribution in which the between study overdispersion depends on the number of experimental units within each HC group (see section 2).
@@ 2 Models for overdispersed binomial data
2.1 Aggregated data
As described above (see fig. 1), HCD is comprised of $h=1,...,H$ historical control groups but further levels of hierarchy can occur within each control group. For example, it is usual in rodent experiments, that within a certain control group, a predefined number of individuals $n_{hi}$ are housed together in one of $i=1,...,I$ cages. Nevertheless, it frequently happens in practice, that dichotomous HCD is documented aggregated on the level of the control groups (omitting the information about the randomization structure within each control group).
A prominent example for this way of reporting is the Historical Control Data Base of the U.S. National Toxicology Program (NTP 2024). In this data base, the total number of rats (or mice) with a certain tumor is presented on control group level (one value per control group) and the information how many rats were housed per cage is only available as meta-data that does not allow to track which rat (or mouse) belonged to a certain cage (as marked by the different shapes in fig. 1). Hence, the aggregation on control group level, leads to the loss of information on the randomization structure within a control group. Consequently, it is not possible to account for potential within control group overdispersion (caused by systematic differences between cages). To account for this type of reporting, also the prediction intervals presented below are based on aggregated data.
If $n_{hi}$ is the number of experimental units per within control group randomization unit (e.g. cages) and $y_{hi}\in\{0,1,...,n_{hi}\}$ is the observed number of successes among them, aggregation on control group level yields
$y_{h}=\sum_{i=1}^{I}y_{hi}$
experimental units with a finding, out of
$n_{h}=\sum_{i=1}^{I}n_{hi}$
experimental units within the whole control group.
2.2 Modeling between-study overdispersion
One possible way to model overdispersion is the beta-binomial distribution in which the binomial proportion for each cluster is derived from a beta distribution
@@ 2.2 Modeling between-study overdispersion
In this setting $E(\pi_{h})=\pi=a/(a+b)$, $E(Y_{h})=n_{h}\pi$, $\pi$ is the overall binomial proportion and $n_{h}$ is the a priori known and fixed total number of experimental units per historical control group. The variance of the beta-binomial distributed random variable is
@@
In this parametrisation $\rho=1/(1+a+b)$ describes the intra-class correlation coefficient (Lui et al. 2000). Please note, that the parameters of the beta distribution are restricted to be $a>0$ and $b>0$. Hence also the intra-class correlation is restricted to be $\rho>0$.
@@
Note, that in the case in which the number of experimental unis within each control group becomes equal ($n_{h}=n_{h^{\prime}}$) also the part of the beta-binomial variance that controls the magnitude of overdispersion per control group ($\big{[}1+(n_{h}-1)\rho\big{]}$ in eq. 3) becomes a constant. Consequently, in this special case, both models are not in contradiction to each other (Menssen and Schaarschmidt 2019) and
$\phi\widehat{=}\big{[}1+(n_{h}-1)\rho\big{]}\text{ if all }n_{h}=n_{h^{\prime}}.$ (5)
@@ 3 Historical control limits for overdispersed binomial data
Since HCL are computed to validate the current control group, the desired HCL should aim to cover the central $100(1-\alpha)$of the underlying distribution and its limits should approach the lower $\alpha/2$ and the upper $21-\alpha/$ 2 quantile, respectively (see fig. 1 of Francq et al. 2019). It it crucial to ensure for this behavior also if the underlying distribution is skewed (Hoffman 2003). Consequently, the desired interval has to ensure for equal tail probabilities
@@
This means that the control limits practically cut off the lower and the upper $100(\alpha/2)$of the underlying distribution, regardless if this distribution is symmetrical or skewed. For overdispersed binomial data, possible skewness depends on three factors:
@@ 3.1 Heuristical historical control limits
they aim to cover all realizations from the underlying distribution and hence, do not aim for a proper approximation of its central $100(1-\alpha)$(but can be rather interpreted as a heuristical 100% prediction interval which does not provide practically useful control limits). Note that the min and max of the observed data provide very poor estimates for the true min and max of the underlying distribution, especially if the number of historical observations is low. This is because values in the center of the distribution have a higher chance to occur, whereas the occurrence of more extreme values in the tails of the distribution are less likely. Furthermore it has to be noted, that this interval is explicitly based on the assumption of constant cluster size $n_{h}=n^{*}\text{ }\forall\text{ }h=1,\ldots H$, because the interpretation of the number of observed events makes only sense in relation to the given cluster size (meaning that 5 out of 50 rats with a tumor have a fundamental different interpretation than 5 out of 100 rats). Hence, a sensible interpretation of the historical range is not possible, if the cluster size differs between the studies.
@@
with $\bar{y}=\frac{\sum_{h}y_{h}}{H}$ and $SD=\sqrt{\frac{\sum_{h}(\bar{y}-y_{h})^{2}}{H-1}}$. This type of HCL is based on a simple normal approximation that lacks an explicit assumption about the mean-variance relationship and hence heuristically allows for overdispersion (Menssen et al. 2024). Similar to HCL that are based on the historical range, the mean $\pm$ $k$ standard deviations are explicitly based on the assumption of constant cluster size $n_{h}=n^{*}\quad\forall\quad h=1,\ldots,H$, because for dichotomous endpoints, the number of observed experimental units with an event is only interpretable in the context of the total number of experimental units in the particular control group. Hence, the application of HCL that are computed by the mean $\pm$ $k$ standard deviations to dichotomous HCD with different sample sizes should strictly be avoided (see section 2 of the supplementary material).
@@ 3.2 Bootstrap calibrated prediction intervals
If between-study overdispersion is modeled based on the quasi-binomial assumption, the corresponding prediction interval is given by
@@
with $\hat{\rho}$ as an estimate for the intra-class correlation coefficient. Details on methodology for the estimation of $\hat{\pi}$, $\hat{\phi}$ and $\hat{\rho}$ are given in section 3.4).
@@
The applied bootstrap calibration is a modified version of the algorithm proposed by Menssen and Schaarschmdit 2022 and has also been successfully applied to overdispersed count data (Menssen et al. 2024). The only difference to Menssen and Schaarschmdit 2022 is, that both interval limits are calibrated individually, but the search for $q^{calib}_{l}$ and $q^{calib}_{u}$ depends on the same bisection procedure using a tolerance $t=0.001$.
Furthermore, the algorithm presented above differs from the calibration procedure used by Menssen and Schaarschmidt 2019 in two important aspects: First, contrary to Menssen and Schaarschmidt 2019 it calibrates both prediction borders separately and hence accounts for equal tail probabilities (see fig. 4 of the supplementary materials). Second, it aims to find substitutes for the complete standard normal quantile used for interval calculation, which in principal also enables simultaneous prediction of the outcome of more than one control group. This is possible if $y^{*}_{b}$ is substituted by a bootstrap sample that mimics the outcome of several control (or treatment) groups.
@@ 3.3.1 Hierarchical modeling
$\pi_{h}\sim\text{Beta}(\mu,\kappa)\quad\forall\;h$
@@ 3.3.2 Bayesian generalized linear mixed model
is drawn from the latter. Similarly as above, a pointwise two-sided prediction interval for one future observation $\tilde{y}^{*}$ is given by
@@ 5 Simulation study
In order to asses the coverage probabilities of the different methods for the calculation of HCL reviewed above, Monte-Carlo simulations were run. In all simulations, the nominal coverage probability was set to $0.951-\alpha=$ 0.95. The parameter combinations used for the simulation were inspired by the real life data shown above. For all simulations the cluster size was fixed ($n_{h}=n^{*}\quad\forall\quad h=1,\ldots,H$) such that the variance-mean relationships of the beta-binomial distribution and the quasi-binomial assumption are not in contradiction to each other.
@@
$\hat{\psi}^{cp}=\frac{\sum_{s=1}^{S}I_{s}}{S}\text{ with}$
@@
Furthermore, the ability of the different historical control limits to ensure for equal tail probabilities was checked. For this purpose the coverage of the simulated lower borders $l_{s}$ and upper borders $u_{s}$ was calculated additionally to the coverage probability for the whole interval. This was done as follows:
$\hat{\psi}^{l}=\frac{\sum_{s=1}^{S}I_{s}}{S}\text{ with}$
$I_{s}=1\text{ if }l_{s}\leq y^{*}_{s},$
$I_{s}=0\text{ if }l_{s}>y^{*}_{s}.$
and
$\hat{\psi}^{u}=\frac{\sum_{s=1}^{S}I_{s}}{S}\text{ with}$ (15)
$I_{s}=1\text{ if }y^{*}_{s}\leq u_{s},$
$I_{s}=0\text{ if }u_{s}>y^{*}_{s}.$
Average interval borders were calculated as the mean of the simulated interval borders
$\bar{l}=\frac{\sum_{s=1}^{S}l_{s}}{S}$
$\bar{u}=\frac{\sum_{s=1}^{S}u_{s}}{S}.$
Each single bootstrap calibrated prediction interval that was calculated in the simulation was based on $B=10000$ bootstrap samples. Each Bayesian prediction interval was calculated based on $C=5000$ MCMC samples.
@@ 5.1 Coverage probabilities for the MNT-inspired simulation setting
As expected, the historical range (fig. 2A) covers the simulated future observations $y^{*}_{s}$ in all the cases, if the amount of historical control groups is high enough and lacks a proper control of the statistical error (here 5 %).
@@
The coverage probabilities of the prediction intervals proposed in sections 3.2 and 3.3 are depicted in Figure 3. In general, both frequentist prediction intervals and the interval based on a Bayesian GLMM yield coverage probabilities relatively close to the nominal level for most settings. All three intervals behave in a comparable fashion, except for settings in which the number of historical control groups is low and / or the amount of overdispersion is high ($H<10$, $\phi=500$). Contrary to this, the prediction interval drawn from a Bayesian hierarchical model, fails to control the statistical error in many of the cases. Furthermore, Bayesian hierarchical models faced serious convergence issues, if the amount of historical studies and the amount of overdispersion was high ($H=100$, $\phi=500$), meaning that the desired interval could not be derived for many of the simulated data sets. Consequently, the coverage probability could not be computed in this three cases (which explains the missing dots for $H=100$ and $\phi=500$ in 3D).
In the (practical) absence of overdispersion ($\phi=1.01$), the beta-binomial prediction interval controls the alpha error best, whereas the prediction intervals based on the quasi-binomial assumption or a Bayesian GLMM tend to behave conservatively. But, in the absence of overdispersion, both intervals approach the nominal coverage probability with an increase of historical control groups. Contrary to this, the prediction interval obtained from Bayesian hierarchical modeling remains conservative in this setting, even for 100 historical studies.

### v4

@@ 1.1 Historical control data (HCD)
Despite the fact, that the HCL based validation of a current control is required by several test guidelines (e.g. OECD 471, OECD 473, OECD 487, OECD 490), most guidelines fail to provide reproducible methodology on how to compute HCL. Hence, many heuristics such as the historical range (the min and the max of the HCD) or the mean $\pm$ 2 standard deviations are used for the practical calculation of HCL, but suffer from severe theoretical drawbacks (Menssen 2023).
1.2 Calculation of historical control limits based on prediction intervals
Contrary to the widely applied heuristics, prediction intervals directly aim to predict one (or more) value(s) from the same data generating process from which the historical data is derived. Hence several authors recommend the use of prediction intervals for the application of HCL (EFSA 2024, Menssen and Schaarschmidt 2019, Menssen and Schaarschmidt 2022, Menssen 2023, Kluxen et al. 2021, Dertinger et al. 2023).
@@ 1.2 Calculation of historical control limits based on prediction intervals
It has to be noted that the prediction intervals presented in several standard textbooks and other scientific publications (e.g. Hahn and Meeker 1991, Hahn et al. 2017, Wang 2008, Tian et al. 2022, Nelson 1982) ground on the assumption that the historical sample is comprised of independent and identically distributed (iid) observations, meaning that possible between-study variation is assumed to be absent. With other words, this simplistic prediction intervals do not account for the hierarchical design of the historical control data, i. e. they do not reflect that certain experimental units belong to a certain historical control group of a certain historical study.
Hence, the restriction to the iid assumption does not make sense in the context of toxicological HCD for which systematic between study variation is a fundamental part of the data generating process, and consequently, is regularly reported in real life data. For example, HC data sets that are prone to systematic between study variation can be found in Tarone 1982, Carlus et al. 2013, Menssen et al. 2024, Levy et al. 2019, Tug et al. 2024, Dertinger et al. 2023 (for further details see section 4 of the supplementary material).
1.3 Dichotomous endpoints
@@
In toxicology, several study types assess dichotomous endpoints: e.g. carcinogenicity studies (rats with vs. rats without a tumor), the micro nucleus test (cells with vs. cells without a micronucleus) or the recently developed liquid microplate fluctuation Ames MPFTM test following Spiliotopoulos et al. 2024 (numbers of wells with vs. number of wells without a color change).
@@
It seems natural to model such data based on the binomial distribution. But, as described before, the hierarchical design of the HCD gives rise for systematic between-study variation. Thus, the variability of the data will exceed the variance that can be modeled by the simple binomial distribution. This effect is called overdispersion (or extra-binomial variation) and can either be modeled based on the quasi-binomial assumption - which relies on the assumption that the between-study overdispersion is constant - or based on the beta-binomial distribution in which the between-study overdispersion depends on the number of experimental units within each HC group (see section 2).
@@ 2 Models for overdispersed binomial data
One possible way to model overdispersion is the beta-binomial distribution in which the binomial proportion for each of the $h=1,\ldots,H$ historical control groups is derived from a beta distribution
@@
In this setting $E(\pi_{h})=\pi=a/(a+b)$, $E(Y_{h})=n_{h}\pi$. Furthermore, $\pi$ is the overall binomial proportion and $n_{h}$ is the a priori known and fixed total number of experimental units per historical control group. The variance of the beta-binomial distributed random variable is
@@
In this parametrisation $\rho=1/(1+a+b)$ describes the intra-class correlation coefficient (Lui et al. 2000). Because the parameters of the beta distribution are restricted to be $a>0$ and $b>0$, also the intra-class correlation is restricted to be $10<\rho<$ 1.
@@
In the case in which the number of experimental unis within each control group becomes equal $n_{h}=n_{1}$, also the part of the beta-binomial variance that controls the magnitude of overdispersion per control group ($\big{[}1+(n_{h}-1)\rho\big{]}$ in eq. 3) becomes a constant. In this condition, the beta-binomial model can be interpreted as a special case of the quasi-binomial model - one in which the magnitude of the intra-class correlation $\rho$ together with the cluster size $n_{h}=n_{1}$ determines the magnitude of the dispersion parameter $\phi$. Consequently, in this special case, both models coincide (Menssen and Schaarschmidt 2019) and
$\phi\widehat{=}\big{[}1+(n_{h}-1)\rho\big{]}\text{ if all }n_{h}=n_{1}.$ (5)
Both approaches are standard for modeling overdispersion (McCullagh and Nelder 1989), but in practice it might sometimes appear by chance, that a data set shows signs of underdispersion (less variability than possible under the simple binomial assumption). From eq. 3 it can be deduced that underdispersion is caused by negative correlations between the experimental units within each cluster (if the intra-class correlation $\rho$ was allowed to be negative). Practically this would mean that the chance to observe a further experimental unit with an event decreases, with the occurrence of an event (e.g. if one rat develops a tumor, the chance that the other rats in the cohort develop a tumor decreases). From a biological point of view, this is heavily implausible. Consequently the quasi-binomial dispersion parameter was restricted to be $\phi\geq 1$ in the following sections of this manuscript.
@@ 3 Historical control limits for overdispersed binomial data
Since HCL are computed to validate the current control group, the desired HCL should aim to cover the central $100(1-\alpha)$of the underlying distribution and its limits should approach the lower $\alpha/2$ and the upper $21-\alpha/$ 2 quantile, respectively (see fig. 1 of Francq et al. 2019). It it crucial to ensure for this behavior also, if the underlying distribution is skewed (Hoffman 2003). Consequently, the desired interval has to ensure for equal tail probabilities
@@
This means that in practice, the control limits should cut off the lower and the upper $100(\alpha/2)$of the underlying distribution, regardless if this distribution is symmetrical or skewed. For overdispersed binomial data, possible skewness depends on three factors:
@@ 3.1 Heuristical historical control limits
they aim to cover all realizations from the underlying distribution and hence, do not aim for a proper approximation of its central $100(1-\alpha)$%. Note that the min and max of the observed data provide very poor estimates for the true min and max of the underlying distribution, especially if the number of historical observations is low. This is because values in the center of the distribution have a higher chance to occur, whereas the occurrence of more extreme values in the tails of the distribution are less likely. Furthermore it has to be noted, that this interval is explicitly based on the assumption of constant cluster size $n_{h}=n^{*}\text{ }\forall\text{ }h=1,\ldots H$, because the interpretation of the number of observed events makes only sense in relation to the given cluster size (meaning that 5 out of 50 rats with a tumor have a fundamental different interpretation than 5 out of 100 rats). Hence, a sensible interpretation of the historical range is not possible, if the cluster size differs between the studies.
@@
with $\bar{y}=\frac{\sum_{h}y_{h}}{H}$ and $SD=\sqrt{\frac{\sum_{h}(\bar{y}-y_{h})^{2}}{H-1}}$. This type of HCL is based on a simple normal approximation that lacks an explicit assumption about the mean-variance relationship and hence heuristically allows for overdispersion (Menssen et al. 2024). Similar to HCL that are based on the historical range, the mean $\pm$ $k$ standard deviations are explicitly based on the assumption of constant cluster size $n_{h}=n^{*}\quad\forall\quad h=1,\ldots,H$, because for dichotomous endpoints, the number of observed experimental units with an event is only interpretable in the context of the total number of experimental units in the particular control group $n_{h}$. Hence, the application of HCL that are computed by the mean $\pm$ $k$ standard deviations to dichotomous HCD with different cluster sizes should strictly be avoided (see section 2 of the supplementary material).
@@ 3.2 Bootstrap calibrated prediction intervals
Following Menssen and Schaarschmidt 2019, a prediction interval which is based on the quasi-binomial assumption is given by
@@
with $\hat{\rho}$ as an estimate for the intra-class correlation coefficient. Details on methodology for the estimation of $\hat{\pi}$, $\hat{\phi}$ and $\hat{\rho}$ are given in section 3.4.
@@
The applied bootstrap calibration is a modified version of the algorithm proposed by Menssen and Schaarschmdit 2022 and has also been applied to calibrate prediction intervals for overdispersed count data (Menssen et al. 2024). The only difference to Menssen and Schaarschmdit 2022 is, that both interval limits are calibrated individually, but the search for $q^{calib}_{l}$ and $q^{calib}_{u}$ depends on the same bisection procedure using a tolerance $t=0.001$.
@@ 3.3.1 Hierarchical modeling
$\pi_{h}\sim\text{Beta}(\mu,\kappa)\quad\forall\;h=1,\ldots,H$
@@ 3.3.2 Bayesian generalized linear mixed model
is drawn from the latter. Similarly as above, a pointwise two-sided prediction interval for one future observation $y^{*}$ is given by
@@ 5 Simulation study
In order to asses the coverage probabilities of the different methods for the calculation of HCL reviewed above, Monte-Carlo simulations were run. In all simulations, the nominal coverage probability was set to $0.951-\alpha=$ 0.95. The parameter combinations used for the simulation were inspired by the real life data shown above. For all simulations the cluster size was fixed ($n_{h}=n^{*}\quad\forall\quad h=1,\ldots,H$) such that the variance-mean relationships of the beta-binomial distribution and the quasi-binomial assumption coincide.
@@
$\hat{\psi}^{cp}=\frac{\sum_{s=1}^{S}I_{s}}{S}\text{ with}$ (15)
@@
Average interval borders were calculated as the mean of the simulated interval borders $\bar{l}=\frac{\sum_{s=1}^{S}l_{s}}{S}$ and $\bar{u}=\frac{\sum_{s=1}^{S}u_{s}}{S}$. Each single bootstrap calibrated prediction interval that was calculated in the simulation was based on $B=10000$ bootstrap samples. Each Bayesian prediction interval was calculated based on $C=5000$ MCMC samples.
@@ 5.1 Coverage probabilities for the MNT-inspired simulation setting
As expected, the historical range (fig. 1A) covers the simulated future observations $y^{*}_{s}$ in all the cases, if the amount of historical control groups is high enough and lacks a proper control of the statistical error (here 5 %). Therefore, the application of the historical range for the calculation of historical control limits should be avoided.
@@
The coverage probabilities of the prediction intervals proposed in sections 3.2 and 3.3 are depicted in Figure 2. In general, both frequentist prediction intervals and the interval based on a Bayesian GLMM yield coverage probabilities relatively close to the nominal level for most settings. All three intervals behave in a comparable fashion, except for settings in which the number of historical control groups is low and / or the amount of overdispersion is high ($H<10$, $\phi=500$). Contrary to this, the prediction interval drawn from a Bayesian hierarchical model, fails to control the statistical error in many of the cases. Furthermore, Bayesian hierarchical models faced serious convergence issues, if the amount of historical studies and the amount of overdispersion was high ($H=100$, $\phi=500$), meaning that the desired interval could not be derived for many of the simulated data sets. Consequently, the coverage probability could not be computed in this three cases (which explains the missing dots for $H=100$ and $\phi=500$ in 2D).
In the (practical) absence of overdispersion ($\phi=1.01$), the beta-binomial prediction interval controls the alpha error best, whereas the prediction intervals based on the quasi-binomial assumption or Bayesian GLMM tend to behave conservatively, but approach the nominal coverage probability with an increase of historical control groups ($H=100$). Contrary to this, the prediction interval obtained from Bayesian hierarchical modeling remains conservative in this setting, even for 100 historical studies.